OpenStudy (anonymous):

Help with math for a MEDAL!!!!!!! The volume of a cone is three hundred cubic kilometers. If the radius of the base is 5 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth.

OpenStudy (anonymous):

please dumb down your answer as much as possible, im terrible at this stuff

OpenStudy (anonymous):

someone please help?

OpenStudy (anonymous):

\[V= \frac{ \pi r^{2} h }{ 3 }\] r= 5 V= 300 Can you plug it in and solve?

OpenStudy (anonymous):

no can you please write it with words so i understand, im not good with symbols

OpenStudy (anonymous):

volume = ( pi times (r square) times h ) / 3

OpenStudy (anonymous):

h= height r= radius pi = 3.14

OpenStudy (anonymous):

well explain what i do step by step in words please

OpenStudy (anonymous):

ok first you put volume = 300 to this equation: volume = ( pi times (r square) times h ) / 3 what do you get?

OpenStudy (anonymous):

replace volume with 300

OpenStudy (anonymous):

would it be 300=(pi) times (r square) times H) /3

OpenStudy (anonymous):

yes now you put r = 5 into this equation : 300=(pi) times (r square) times H) /3 what do you get?

OpenStudy (anonymous):

300=(pi) times (r 5) times H) /3

OpenStudy (anonymous):

no, it's would it be 300=(pi) times (5 square) times H) /3

OpenStudy (anonymous):

now pi = 3.14 so replace pi with 3.14 in this equation: 300=(pi) times (5 square) times H) /3

OpenStudy (anonymous):

ok well do i put 300=(pi) times (5 square) times H) /3 first on my exam or what?

OpenStudy (anonymous):

is the exam asking you to show work?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then first, put down the equation \[V= \frac{ \pi r^{2} h }{ 3 }\] then write that the equation you just brought up down.

OpenStudy (anonymous):

well just show me how to write the whole thing, like how would i type it?

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

do you know how to type these symbols? V= (pi r^2 (h)) / 3

OpenStudy (anonymous):

yes but are you sure this is right? i just want the answer typed out in words, no symbols like "Pi 2 r gjnjihjtrienjtyh+ + = 1/3 69" I want it typed out in words, the process of how to do this

OpenStudy (anonymous):

like i want you to be like "than you multiply this and that divide this"

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

anyone there?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

can someone just give me the answer without symbols and all this confusing stuff? can someone write it out in WORDS you will get a medal!

OpenStudy (anonymous):

HELP!!!!

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

\[V= \frac{ pir^{2}h }{ 3 }\] \[300 = \frac{ \pi (25) h }{ 3 }\] \[\frac{ 300(3) }{ 25 \pi } = h\] \[h= \frac{ 36 }{ \pi } = 11. 46\]

OpenStudy (anonymous):

but you need to know how to do this on your own. Getting the right answer is not as important as understasnding how to do this.

OpenStudy (anonymous):

i know, what im asking for you to do is explain this to me in words not symbols

OpenStudy (anonymous):

say stuff like "then you multiply this then that, add that, subtract that"

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

@study100 r u there?

OpenStudy (anonymous):

or can you only explain in symbols, is it not possible to write it in words, if it is, i understand but from the equations you are shwoing me, i dont know if im suppose to multiply something or divide something

OpenStudy (anonymous):

volume = pi r squared times h all over 3 three hundred = thee point four times twenty five times h all over three three hundred times three = three point four times twenty five times h

OpenStudy (anonymous):

now you divide three point four to both sides

OpenStudy (anonymous):

ok now that makes less sense just explain to me with symbols and words like show me the "3/4/69 8 5 pi r s ZZZ" stuff and explain to me what to do, because i just want to get this questions done, i want to know this! please

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

ok first of all, do you know how to plug in formulas?

OpenStudy (anonymous):

do you know the area of the circle? 3. do you know what pi is?

OpenStudy (anonymous):

do you know about radius/diameter/ volume of a cone? What makes up a cone?

OpenStudy (anonymous):

pi is 3.14

OpenStudy (anonymous):

ok good, what is the formula for area of circle?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

isnt it 1/3

OpenStudy (anonymous):

no you've got a long a way to go. Let's start with basics first, how many degrees is in a circle?

OpenStudy (anonymous):

idk cant you just literally write the answer, im never gonna use this in life, you can have symbols just in between symbols and all that stuff write "and then i did" or something

OpenStudy (anonymous):

@study100